28 marzo 2014

Observer: F. Cusano, R. Carini, M. Faccini
Telescope Operator: G. Beckett
Support Astronomer: O. Kuhn (on call)
Instrument: MODS

Summary

Late afternoon the humidity is 102%, outside temperature -2.5 C, dew point temperature -2.2 C, wind around 14 m/s.
Due to the dew point limit, we can open the dome, and the temperatures are : chamber=3.48 C, glass = -0.47 C
Heavy clouds just on top of the telescope.
the situation go better and we open the dome: temperatures of the glasses are: -0.4 C on the left and -0.8 C on the right, the external is -4 C and the chamber is -3 C.
Problems with the temperature gradient on the mirrors, the focus is not very good.
During the transit we had problems with the blue shutter. It remains stuck opened and we need to stop the ob and operate with the SA every one/two images.
During the transit the seeing value oscillate between 1.4 and 2.9.
We lost about 17 minutes of images during the last part of the transit.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Bellazzini_B HVC204.88+44.86+147 g,r 11m COMPLETED
Bellazzini_A HVC205.28+18.70+150 g,r 11m COMPLETED
Bellazzini_K HVC196.09+24.74+166 g,r 11m to check the last image
Clementini_DDT_2 IZw18 B,r 10m COMPLETED
Nascimbeni_33 gj436 U_spec, F972N20 2.5h COMPLETED , problem with blue shutter in the last hour
Bellazzini_S HVC27.86+38.25+124 g,r 11m COMPLETED
Standard SA101 B, V, g; I, r, r    
Standard SA107 B, V, g; I, r, r   problem with the blue shutter

Details

02:00 The TO opens the dome

02:01 we star sky flat
  • flat in U_spec and F972N20 first good file lbcb.20130328.020342- lbrcr020339, the counts of the F972N20 filter are low
  • flat g and I lbcb021128-lbcr021125 good
  • flat V and r lbcb022049-lbcr022034 counts too low, to repeat
  • 02:24 finish with sky flat

02:42 slew to the standard SA101

02:45 run dofpia

02:50 the temperatures of the glasses are: -0.4 C on the left and -0.8 C on the right, the external is -4 C and the chamber is -3 C.
  • There are problems with the temperature gradient on the mirrors, the focus is not very well, overall on the left side.
  • S.A. suggest to stay on LBC anyway because of the transit

3:00 trying again dofpia

3.10 standard field SA101, run SA101_BVg_Ir.ob

3.15 slew to Bellazzini, run 6_B_copointing.ob
  • primary panic, have to restart the mirror

3:40 slew to zenith to restore the primary mirror

3.46 run again 6_B_copointing.ob
  • dofpia
  • no pupils found in the red channel, start again dofpia, there is a big shift between the blue and red camera so we run first the copointing

3:52 lbcrangebal

3:57 slew to science, run 6_B.ob
  • dofpia

4:06 we still see a big shift between the blue and red camera, we run lbcrangebal again
  • lbcrangebal
  • preset cancel!
  • lbcrangebal, again preset cancel
  • problem in sending the preset of the lbcrangebal, the TO reset the correction
  • lbcrangebal

4:36 run 6_B.ob, Bellazzini
  • seeing 1.1 in the blue chanel, seeing 0.9 in the red chanel

04:52 slew to Bellazzini_A
  • dofpia
  • stop dofpia, sending manual corrections
  • send dofpia again

5:01 manual correction by the TO due to coma
  • dofpia
  • manual correction by the TO
  • dofpia

05:17 run 6_A.ob, Bellazzini

05:32 slew to 6_K.ob, Bellazzini
  • dofpia

05:38 run 6_K.ob, Bellazzini
  • problem in the last images

05:55 slew to Clementini

05:56 blue shutter didn't close. error in GUI:
  • error messages "5:50:36 shutter did not open" "5:50:37 camera system encountered a problem: shutter hardware failure"

6:00 problem resolved by SA
  • dofpia

6:12 stopped dofpia, it didn't converge

6:17 dofpia sent after the manual spherical changes
  • still problems, it didn't converge

6:32 manual corrections by TO

6:42 dofpia

6:47 Run IZw18_2.ob
  • seeing 0.7

7:00 slew to Nascimbeni_33, send to copointing RA11h42m_DEC26d42m.ob
  • clear the forces manually
  • dofpia

7:17 the blue chanel did not converge, stop the dofpia

7.18 send to copointing RA11h42m_DEC26d42m.ob
  • lbcrangebal

7:25 start gj436_test.ob
  • dofpia
  • defocus

7.32 run gj436_test.ob ==lbcb.20140328.
  • the values in the red side are too high, in the blue side are too low
  • change the value of the defocus manually

7.38 run again gj436_test.ob
  • some pixel in the red side have value higher than 35000
  • defocus only the red side * run gj436_test.ob

7:51 run again gj436_test.ob

7:57 run again gj436_test.ob

7.59 send the first ob gj436_1.ob

8:03 stop the ob to reduce the counts of the reference star #4

8:04 start again ob gj436_1.ob== lbc.20140328.080426 lbcb.20140328.080438==

8:34 run gj436_2.ob

8:50 the PI is contacted to discuss a possible problem of proximity to the CCD limit

9.06 run gj436_2_bis.ob in which is changed the number of row of the CCD read (set Y=1900 instead of 2200)

9.06 run gj436_3_bis.ob in which is changed the number of row of the CCD read (set Y=1900 instead of 2200)
  • we have changed also gj436_4 and gj436_5 in the same way

9.36 run gj436_4_bis.ob

9:48 stop the gj436_4_bis.ob because the shutter closed

9.49 problem solved, run again gj436_4_bis.ob

9:53 error message (shutter opened), stop the ob

9:54 reload gj436_4_bis.ob
  • run again gj436_4_bis.ob
  • shutter stuck, stop the ob

9:56 reload the gj436_4_bis.ob
  • run gj436_4_bis.ob
  • shutter stuck, stop the ob

10:00 reload the gj436_4_bis.ob
  • run gj436_4_bis.ob

10:05 the blue shutter is locked, we stop the ob
  • run again gj436_4_bis.ob

10:09 error message again (shutter stuck), stop the ob
  • run gj436_5_bis.ob

10:16 shutter stuck
  • run gj436_5_bis.ob

10:20 shutter stuck
  • run gj436_5_bis.ob

10:40 we do the requested bias for Nascimbeni transit and the bias for the modified ob-script

10:49 slew to Bellazzini 6S
  • dofpia

11:02 error message in the blue channel, stop dofpia and run again
  • dofpia, /x2, /backout
  • dofpia, /x2

11.12 error message shutter blue,
  • stop dofpia
  • dofpia, /x2, /first
  • copointing

11:23 run 6_S.ob, Bellazzini
  • 11:29 blue shutter stuck
  • 11:45 shutter failure in the blue
  • red side images completed, blue side images 1x300s and 1x20s done
  • 11:19 run again 6_S.ob, blue images have problem

11:52 run again 6_S.ob, Bellazzini

12:00 slew to standard field SA107
  • run SA107_BVg_rI.ob
  • blue shutter stuck, we are trying to recover
  • a lot of problem with the shutter, the images need to be checked

12:17 starting skyflat
  • g,I flat
  • shutter stuck......
  • no flat

12:39 end of the night, the night of the blue shutter
Topic revision: r8 - 29 Mar 2014, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback